# Weak interaction axial form factors of the octet baryons in nuclear medium

### v1

@@ III.2 Meson cloud contribution
We consider now the effect of the meson cloud. When the baryon wave function has contributions from valence quarks and from the meson cloud we can consider the following decomposition
@@
where $Z_{B}$ represent normalization factor determined in the study of the baryon $B$ electromagnetic factors [28]. In the transition $n\to p$, we can replace $\sqrt{Z_{B^{\prime}}Z_{B}}\to Z_{N}$.
From the comparison with Eq. (4) we can write $G_{A}^{\rm B}=\sqrt{Z_{B^{\prime}}Z_{B}}\tilde{G}_{A}^{\rm B}$ and $G_{A}^{\rm MC}=\sqrt{Z_{B^{\prime}}Z_{B}}\tilde{G}_{A}^{\rm MC}$.
@@
where $\mu=m_{\pi}$ (pion mass) for $\Delta I=1$ transitions and $\mu=m_{K}$ (kaon mass) for $\Delta S=1$ transitions, as discussed in Ref. [8].
To simplify the terminology, we call the tilde quantities ($\tilde{G}_{A}^{\rm B}$, $\tilde{G}_{P}^{\rm B}$, $\tilde{G}_{A}^{\rm MC}$, $\tilde{G}_{P}^{\rm pole}$ and $\tilde{G}_{A}^{\rm MC}$, bare, MC or pole) as relative contributions to the form factors, and the relations corrected by the wave function normalization factors as the effective contributions to the form factors. Notice that the final result require the factor $\sqrt{Z_{B^{\prime}}Z_{B}}$, while the bare contributions are naturally independent of these factors. At large $Q^{2}$ the meson cloud contribution is suppressed and $G_{A}\simeq\sqrt{Z_{B^{\prime}}Z_{B}}G_{A}^{\rm B}$. In the present work, we assume that the dominant contribution is the pion cloud and use the relations from Ref. [28] to determine $Z_{B}$.
@@
We can now discuss the function $\tilde{G}_{A}^{\rm MC}$ which determine the meson cloud contribution to the form factors $G_{A}$. For $\tilde{G}_{A}^{\rm MC}$, we consider an effective parametrization based on the meson-baryon SU(3) coupling, with two independent functions $F$ and $D$ of $Q^{2}$ [5, 8]. The value of the functions for $Q^{2}=0$ can be determined by the results for $G_{A}(0)$ for different known transitions ($n\to p$, $\Lambda\to p$, $\Sigma^{-}\to n$, $\Xi^{-}\to\Lambda$ and $\Xi^{0}\to\Sigma^{+}$). The explicit values are
$F_{0}\equiv F(0)=0.1775,\hskip 11.38092ptD_{0}\equiv D(0)=0.4284.$ (27)
@@
In the extension from the nucleon to the octet baryon cases we replace $F$ and $D$ by functions of $Q^{2}$, according to the $Q^{2}$-dependence observed for the meson cloud contribution to the $n\to p$ transition, as described next.
The present calibration differs from meson-baryon SU(3) model which combines the bare and meson cloud contributions by the values of $F$ and $D$ [5]. Thus, we consider here a modified version, where, only the meson cloud contributions are described by the meson-baryon SU(3) symmetry. The bare contributions is determined by our approximated SU(6) quark model, where the breaking is included at the level of the radial wave functions with parameters depending on the number of strange quarks, and by the values of the baryon masses.
@@
$\Delta I=1\hskip 20.00003pt$       $n\to p$ $F+D$ $\frac{F_{0}+D_{0}}{H_{0}}$
      $\Sigma^{\pm}\to\Lambda$$\pm\sqrt{\frac{2}{3}}D$$\pm\sqrt{\frac{2}{3}}\frac{D_{0}}{H_{0}}$
        $\Sigma^{-}\to\Sigma^{0}$ $\sqrt{2}F$ $\sqrt{2}\frac{F_{0}}{H_{0}}$
        $\Sigma^{0}\to\Sigma^{+}$ $\sqrt{2}F$ $\sqrt{2}\frac{F_{0}}{H_{0}}$
      $\Xi^{-}\to\Xi^{0}$ $F-D$ $\frac{F_{0}-D_{0}}{H_{0}}$
$\Delta S=1\hskip 20.00003pt$       $\Lambda\to p$ $-\sqrt{\frac{3}{2}}\left(F+\frac{1}{3}D\right)$ $-\sqrt{\frac{3}{2}}\frac{F_{0}+D_{0}/3}{H_{0}}$
      $\Sigma^{-}\to n$ $-F+D$ $-\frac{F_{0}-D_{0}}{H_{0}}$
     $\Sigma^{0}\to p$ $-\frac{1}{\sqrt{2}}(F-D)$ $-\frac{1}{\sqrt{2}}\frac{F_{0}-D_{0}}{H_{0}}$
      $\Xi^{-}\to\Lambda$ $-\sqrt{\frac{3}{2}}\left(F-\frac{1}{3}D\right)$ $-\sqrt{\frac{3}{2}}\frac{F_{0}-D_{0}/3}{H_{0}}$
       $\Xi^{-}\to\Sigma^{0}$ $\frac{1}{\sqrt{2}}(F+D)$ $\frac{1}{\sqrt{2}}\frac{F_{0}+D_{0}}{H_{0}}$
       $\Xi^{0}\to\Sigma^{+}$ $F+D$ $\frac{F_{0}+D_{0}}{H_{0}}$
$\,\Delta I=0,\hskip 20.00003pt$ $N\to N$ $(F+D)\tau_{3}$ $\frac{F_{0}+D_{0}}{H_{0}}\tau_{3}$
$\Delta S=0\hskip 20.00003pt$ $\Sigma\to\Sigma$ $2FI_{\Sigma}$ $2\frac{F_{0}}{H_{0}}I_{\Sigma}$
$\Xi\to\Xi$ $(F-D)\tau_{3}$ $\frac{F_{0}-D_{0}}{H_{0}}\tau_{3}$
Table 4: Octet baryon axial-vector form factors $G_{A}(Q^{2})$, expressed in terms of $F$ and $D$ of SU(3) scheme [5, 45, 46]. $F_{0}$ and $D_{0}$ are defined by Eq. (27) and $H_{0}=F_{0}+D_{0}$. The last rows refer to the neutral currents. The contributions to $\Lambda$ and $\Sigma^{0}$ are zero [8].
The meson cloud contribution for the nucleon axial-vector form factor is determined by the fit to the nucleon data based on the results for $G_{A}$ estimated by the lattice QCD data [38]
@@
where $G_{A}^{{\rm MC}0}=0.6059$ and $\Lambda=1.05$ GeV. In the notation of Table 4: $G_{A}^{{\rm MC}0}=H_{0}$.
@@
where $\tilde{G}_{A,N}^{\rm MC}$ represents the parametrization of the nucleon’s meson cloud contribution given by Eq. (28) and $\eta_{BB^{\prime}}$ is the coefficient associated with the SU(3) symmetry displayed in the last column of Table 4. Notice that the coefficient $\eta_{BB^{\prime}}$ is a constant defined by the values of $F_{0}$ and $D_{0}$ and the $Q^{2}$-dependence comes exclusively from the function $\tilde{G}_{A,N}^{\rm MC}$.
$G_{A}^{\rm MC}(Q^{2})=\sqrt{Z_{B^{\prime}}Z_{B}}\;\tilde{G}_{A}^{\rm MC}(Q^{2}).$ (30)
In Table 4 we represent some factors $\eta_{BB^{\prime}}$ by using the factor $(F_{0}+D_{0})/H_{0}$ when we could have used the unity. This representation is intentional in order to prepare the generalization for a case where the meson-baryon coupling may change, as discussed in the next section (extension to the nuclear medium). In that case we consider that the functions $\tilde{G}_{A}^{\rm MC}$ change with the (squared) coupling constants. This can be done assuming that $F_{0}$ and $D_{0}$ include the dependence on the coupling constants, while $H_{0}$ is constant determined numerically and independent of the coupling constants.
Like the meson cloud contributions, the bare contributions are broken by the normalization associated with the baryon wave functions (factors $\sqrt{Z_{B}}$). In these conditions the meson cloud contributions for $G_{A}$ for the transitions $n\to p$ and $\Xi^{0}\to\Sigma^{+}$, determined by $F+D$, are not the same since the suppression associated with the factors $\sqrt{Z_{B}}$ is stronger for the nucleon (smaller $\sqrt{Z_{B}}$).
In the present work we consider a modification of the parametrization from Ref. [8], in order to correct small numerical errors in the calculation of $Z_{\Xi}$ and $Z_{\Sigma}$. We use this opportunity to calibrate the parametrization for $F_{0}$ and $D_{0}$ since the previous estimates of the $\Xi$ decays were slightly overestimated. The new parameters are displayed in Eq. (27). Compared with Ref. [8] the corrections are very small (less than 0.5%), but we consider the new values for consistency. The main difference between the calculations are due to the factor $Z_{\Xi}$, underestimated about 10%. The numerical results for the axial-vector form factors in vacuum are presented and discussed in Sec. V.1.
@@ III.3 Combination of bare and meson cloud
For the extension of the formalism to the nuclear medium it is convenient to relate the effective coefficients $F$ and $D$ used in Eqs. (29) and (30) to the explicit meson cloud mechanisms and the meson-baryon couplings. From the analysis of the meson cloud diagrams depicted in Fig. 1, one can conclude that the meson cloud contribution is proportional to $g_{MBB_{1}}g_{MB_{1}B^{\prime}}$ where $M$ is the dressing meson ($\pi$, $K$, …) and $B_{1}$ is an intermediate baryon state (octet or decuplet). Since, in the effective SU(3) meson-baryon interactions we can write all $g_{MBB^{\prime}}$ in terms of a coefficient times $g_{\pi NN}$, we conclude that
@@ IV Extension of the model to symmetric nuclear medium
We discuss now the extension of the model for the vacuum described in the previous section to a symmetric nuclear matter. The extension is based on the modification of the properties of the hadrons and the meson-baryon coupling constants to the nuclear matter. These properties can be estimated by the quark-meson coupling model applied to the symmetric nuclear matter [23, 24].
In the extension of our model to the nuclear matter, we consider the values of the hadron masses (baryons and mesons) $m_{h}$ from the vacuum to the medium $m_{h}^{\ast}$, as well as the modifications of the coupling constants $g_{MBB^{\prime}}$ in vacuum to $g_{MBB^{\prime}}^{\ast}$ in medium. The consequence of this extension is that the bare contributions to the axial form factors are modified, since we replace $m_{\rho}$, $m_{\phi}$ and $M_{h}=2M_{N}$ respectively by $m_{\rho}^{\ast}$, $m_{\phi}^{\ast}$ and $M_{h}^{\ast}=2M_{N}^{\ast}$ in the quark axial current and the radial wave functions of the baryon $B$ are modified replacing the baryon mass $M_{B}$ by the effective baryon mass in medium $M_{B}^{\ast}$. As for the meson cloud contributions, they change as a consequence of the modification of the coupling $g_{MBB^{\prime}}$ to $g_{MBB^{\prime}}^{\ast}$.
$\rho/\rho_{0}$ $M_{N}$ $m_{\rho}$ $g_{\pi NN}^{\ast}/g_{\pi NN}$ $g_{A}^{N\ast}/g_{A}^{N}$ $G_{A}^{N}(0)$ $n_{P}$
@@
In addition to the bare contributions ($\tilde{G}_{A}^{\rm B}$ replaced by $\tilde{G}_{A}^{{\rm B}*}$), the effect of the nuclear medium acts in two levels: on the meson cloud contribution ($\tilde{G}_{A}^{\rm MC}$ replaced by $\tilde{G}_{A}^{{\rm MC}*}$), and on the wave function normalization factors $Z_{B}$ and $Z_{B^{\prime}}$ (global modification of $\sqrt{Z_{B}Z_{B^{\prime}}}$ to $\sqrt{Z_{B}^{\ast}Z_{B^{\prime}}^{\ast}}$).
We discuss the modifications on the function $\tilde{G}_{A}^{{\rm MC}}$ first. From the discussion from Eq. (29) relative to the meson cloud contribution to a generic $B\to B^{\prime}$ transition, we conclude that $\tilde{G}_{A}^{\rm MC}$ include linear combinations of $F$ and $D$, which are proportional to the product of two coupling constants $g_{\pi BB^{\prime}}^{2}$, since these include diagrams with two pion-baryon couplings (see Fig. 1). The combination of coupling constants can be expressed in terms of $g_{\pi NN}^{2}$, as discussed already. The conclusion is then that $\tilde{G}_{A}^{{\rm MC}}\propto g_{\pi NN}^{2}$ and we can write
@@
In the extension to the nuclear medium we assume, for simplicity, that medium effects associated with the meson cloud contributions are taken into account using the modified coupling constants and normalization factors. Thus, in medium $\tilde{G}_{A,N}^{\rm MC}$ is determined by Eq. (28) where the cutoff $\Lambda$ value determined by nucleon data in vacuum.

### v2

@@ III.2 Meson cloud contribution
We consider now the effect of the meson cloud. Following Ref. [7], we consider a phenomenological description of the meson cloud effects. The use of a phenomenological description was motivated by the fact that the magnitude of the meson cloud contributions is expected to be smaller than the valence quark contributions [7, 9, 28], and by simplicity.
When the baryon wave function has contributions from valence quarks and from the meson cloud we can consider the following decomposition
@@
where $Z_{B}$ represents the normalization factor determined in the study of the baryon $B$ electromagnetic factors [28]. The bare contribution is associated with the short-range interactions, and the meson cloud contribution is associated with the long-range interactions.
From the comparison with Eqs. (4) and (5) we can write $G_{A}^{\rm B}=\sqrt{Z_{B^{\prime}}Z_{B}}\tilde{G}_{A}^{\rm B}$ and $G_{A}^{\rm MC}=\sqrt{Z_{B^{\prime}}Z_{B}}\tilde{G}_{A}^{\rm MC}$. In the transition $n\to p$, we can replace $\sqrt{Z_{B^{\prime}}Z_{B}}\to Z_{N}$.
As mentioned already, the decomposition from Eq. (23) into a bare and meson cloud contributions is model dependent [3, 121]. This model dependence is a consequence of the ambiguities on the calibration of the background and in the identification of the bare states [117, 118, 119, 122, 121]. The impact of the model dependence can be reduced by matching the valence quark contributions of our model with the results from lattice QCD simulations for large pion masses, as explained next. Although quenched and unquenched lattice QCD simulations include some meson cloud effects, those effects are reduced for larger pion masses [121, 123].
@@
where $\mu=m_{\pi}$ (pion mass) for $|\Delta I|=1$ transitions and $\mu=m_{K}$ (kaon mass) for $|\Delta S|=1$ transitions, as discussed in Ref. [7].
To simplify the terminology, we call the tilde quantities ($\tilde{G}_{A}^{\rm B}$, $\tilde{G}_{P}^{\rm B}$, $\tilde{G}_{A}^{\rm MC}$, $\tilde{G}_{P}^{\rm pole}$ and $\tilde{G}_{A}^{\rm MC}$, bare, MC or pole) as relative contributions to the form factors, and the relations corrected by the wave function normalization factors as the effective contributions to the form factors. Notice that the final result requires the factor $\sqrt{Z_{B^{\prime}}Z_{B}}$, while the bare contributions are naturally independent of these factors. At large $Q^{2}$ the meson cloud contribution is suppressed and $G_{A}\simeq\sqrt{Z_{B^{\prime}}Z_{B}}\tilde{G}_{A}^{\rm B}$. In the present work, we assume that the dominant contribution is the pion cloud and use the relations from Ref. [28] to determine $Z_{B}$.
@@
We can now discuss the function $\tilde{G}_{A}^{\rm MC}$ which determines the meson cloud contribution to the form factors $G_{A}$. For $\tilde{G}_{A}^{\rm MC}$, we consider an effective parametrization based on an SU(3) meson-baryon coupling, with two independent functions $F^{\prime}$ and $D^{\prime}$ of $Q^{2}$ [5, 7]. We use primes to emphasize the difference to the standard SU(3) meson-baryon model, where the functions parametrize the combination of valence and meson cloud contributions. The value of the functions for $Q^{2}=0$ can be determined by the results for $G_{A}(0)$ for different known transitions ($n\to p$, $\Lambda\to p$, $\Sigma^{-}\to n$, $\Xi^{-}\to\Lambda$ and $\Xi^{0}\to\Sigma^{+}$). The explicit values are
$F_{0}\equiv F^{\prime}(0)=0.1775,\hskip 11.38092ptD_{0}\equiv D^{\prime}(0)=0.4284.$ (27)
@@
In the extension from the nucleon to the octet baryon cases we replace $F^{\prime}$ and $D^{\prime}$ by functions of $Q^{2}$, according to the $Q^{2}$ dependence observed for the meson cloud contribution to the $n\to p$ transition, as described next.
We emphasize that we consider here a modified version, where, only the meson cloud contributions are described by a meson-baryon SU(3) symmetry model. The bare contributions are determined by our approximated SU(6) quark model, where the breaking is included at the level of the radial wave functions with parameters depending on the number of strange quarks, and by the values of the baryon masses.
@@
$|\Delta I|=1\hskip 20.00003pt$       $n\to p$ $F^{\prime}+D^{\prime}$ $\frac{F_{0}+D_{0}}{H_{0}}$
      $\Sigma^{\pm}\to\Lambda$$\pm\sqrt{\frac{2}{3}}D^{\prime}$$\pm\sqrt{\frac{2}{3}}\frac{D_{0}}{H_{0}}$
        $\Sigma^{-}\to\Sigma^{0}$ $\sqrt{2}F^{\prime}$ $\sqrt{2}\frac{F_{0}}{H_{0}}$
        $\Sigma^{0}\to\Sigma^{+}$ $\sqrt{2}F^{\prime}$ $\sqrt{2}\frac{F_{0}}{H_{0}}$
      $\Xi^{-}\to\Xi^{0}$ $F^{\prime}-D^{\prime}$ $\frac{F_{0}-D_{0}}{H_{0}}$
$|\Delta S|=1\hskip 20.00003pt$       $\Lambda\to p$ $-\sqrt{\frac{3}{2}}\left(F^{\prime}+\frac{1}{3}D^{\prime}\right)$ $-\sqrt{\frac{3}{2}}\frac{F_{0}+D_{0}/3}{H_{0}}$
      $\Sigma^{-}\to n$ $-F^{\prime}+D^{\prime}$ $-\frac{F_{0}-D_{0}}{H_{0}}$
     $\Sigma^{0}\to p$ $-\frac{1}{\sqrt{2}}(F^{\prime}-D^{\prime})$ $-\frac{1}{\sqrt{2}}\frac{F_{0}-D_{0}}{H_{0}}$
      $\Xi^{-}\to\Lambda$ $-\sqrt{\frac{3}{2}}\left(F^{\prime}-\frac{1}{3}D^{\prime}\right)$ $-\sqrt{\frac{3}{2}}\frac{F_{0}-D_{0}/3}{H_{0}}$
       $\Xi^{-}\to\Sigma^{0}$ $\frac{1}{\sqrt{2}}(F^{\prime}+D^{\prime})$ $\frac{1}{\sqrt{2}}\frac{F_{0}+D_{0}}{H_{0}}$
       $\Xi^{0}\to\Sigma^{+}$ $F^{\prime}+D^{\prime}$ $\frac{F_{0}+D_{0}}{H_{0}}$
$\,\Delta I=0,\hskip 20.00003pt$ $N\to N$ $(F^{\prime}+D^{\prime})\tau_{3}$ $\frac{F_{0}+D_{0}}{H_{0}}\tau_{3}$
$\Delta S=0\hskip 20.00003pt$ $\Sigma\to\Sigma$ $2F^{\prime}I_{\Sigma}$ $2\frac{F_{0}}{H_{0}}I_{\Sigma}$
$\Xi\to\Xi$ $(F^{\prime}-D^{\prime})\tau_{3}$ $\frac{F_{0}-D_{0}}{H_{0}}\tau_{3}$
Table 4: Octet baryon axial-vector form factors $G_{A}(Q^{2})$, expressed in terms of $F^{\prime}$ and $D^{\prime}$ of SU(3) scheme [5, 45, 46]. $F_{0}$ and $D_{0}$ are defined by Eq. (27) and $H_{0}=F_{0}+D_{0}$. The last rows refer to the neutral currents. The contributions to $\Lambda$ and $\Sigma^{0}$ are zero [7].
The meson cloud contribution for the nucleon axial-vector form factor is determined by the fit to the nucleon data based on the results for $G_{A}$ estimated by the lattice QCD data
@@
where $G_{A}^{{\rm MC}0}=0.6059$ and $\Lambda=1.05$ GeV. In the notation of Table 4, $G_{A}^{{\rm MC}0}=H_{0}$. In Ref. [7], we demonstrated the advantage of using a SU(3) meson-baryon parametrization to describe the meson cloud component of the model.
where $\tilde{G}_{A,N}^{\rm MC}$ represents the parametrization of the nucleon’s meson cloud contribution given by Eq. (28) and $\eta_{BB^{\prime}}$ is the coefficient associated with the SU(3) symmetry displayed in the last column in Table 4. Notice that the coefficient $\eta_{BB^{\prime}}$ is a constant defined by the values of $F_{0}$ and $D_{0}$ and the $Q^{2}$ dependence comes exclusively from the function $\tilde{G}_{A,N}^{\rm MC}$.
$G_{A}^{\rm MC}(Q^{2})=\sqrt{Z_{B^{\prime}}Z_{B}}\;\tilde{G}_{A}^{\rm MC}(Q^{2}).$ (30)
In Table 4 we represent some factors $\eta_{BB^{\prime}}$ by using the factor $(F_{0}+D_{0})/H_{0}$ when we could have used unity. This representation is intentional in order to prepare the generalization for a case where the meson-baryon coupling may change, as discussed in the next section (extension to the nuclear medium). In that case we consider that the functions $\tilde{G}_{A}^{\rm MC}$ change with the (squared) coupling constants. This can be done by assuming that $F_{0}$ and $D_{0}$ include the dependence on the coupling constants, while $H_{0}$ is constant determined numerically and independent of the coupling constants.
Like the meson cloud contributions, the bare contributions are broken by the normalization associated with the baryon wave functions (factors $\sqrt{Z_{B}}$). In these conditions the meson cloud contributions for $G_{A}$ for the transitions $n\to p$ and $\Xi^{0}\to\Sigma^{+}$, determined by $F^{\prime}+D^{\prime}$, are not the same since the suppression associated with the factors $\sqrt{Z_{B}}$ is stronger for the nucleon (smaller $\sqrt{Z_{B}}$).
@@
In the present work we consider a modification of the parametrization from Ref. [7], in order to correct small numerical errors in the calculation of $Z_{\Xi}$ and $Z_{\Sigma}$. We use this opportunity to calibrate the parametrization for $F_{0}$ and $D_{0}$ since the previous estimates of the $\Xi$ decays were slightly overestimated. The new parameters are displayed in Eq. (27). Compared with Ref. [7] the corrections are very small (less than 0.5%), but we consider the new values for consistency. The main difference between the calculations is due to the factor $Z_{\Xi}$, underestimated about 10%. The numerical results for the axial-vector form factors in vacuum are presented and discussed in Sec. V.1.
@@ III.3 Combination of bare and meson cloud
For the extension of the formalism to the nuclear medium it is convenient to relate the effective coefficients $F^{\prime}$ and $D^{\prime}$ used in Eqs. (29) and (30) to the explicit meson cloud mechanisms and the meson-baryon couplings. From the analysis of the meson cloud diagrams depicted in Fig. 1, one can conclude that the meson cloud contribution is proportional to $g_{MBB_{1}}g_{MB_{1}B^{\prime}}$ where $M$ is the dressing meson ($\pi$, $K$, …) and $B_{1}$ is an intermediate baryon state (octet or decuplet). Since, in the effective SU(3) meson-baryon interactions we can write all $g_{MBB^{\prime}}$ in terms of a coefficient times $g_{\pi NN}$, we conclude that
@@ IV Extension of the model to symmetric nuclear medium
We discuss now the extension of the model for the vacuum described in the previous section to symmetric nuclear matter. The extension is based on the modification of the properties of the hadrons and the meson-baryon coupling constants to the nuclear matter. These properties can be estimated by the QMC model applied to the symmetric nuclear matter [23, 24].
The standard QMC model which we use is based on the MIT bag model, to be able to describe the symmetric nuclear matter reasonably well by using the $\sigma$ and $\omega$ mean fields self-consistently. The self-consistent exchange of the Lorentz-scalar-isoscalar $\sigma$ and Lorentz-vector-isoscalar $\omega$ mean fields among the light quarks in nucleons provides a good description of the symmetric nuclear matter, with the incompressibility value of about 300 MeV (within the extracted empirical range), without introducing complicated nonlinear higher-order meson interactions. Based on this rather simple treatment with very less number of parameters, the QMC model has been applied successfully for many nuclear phenomena (see e.g., Refs. [23, 24]). The in-medium effective masses of hadrons are controlled by the values of the self-consistently obtained $\sigma$ mean field. The coupling constants in medium are estimated using the Goldberger–Treimann relation and the medium values of the baryon masses, baryon axial-couplings and pion decay constant $f_{\pi}$ in medium.
In the extension of our model to nuclear matter, we consider the values of the hadron masses (baryons and mesons) $m_{h}$ from the vacuum to the medium $m_{h}^{\ast}$, the vector potentials, as well as the modifications of the coupling constants $g_{MBB^{\prime}}$ in vacuum to $g_{MBB^{\prime}}^{\ast}$ in medium. The consequence of this extension is that the bare contributions to the axial form factors are modified, since we replace $m_{\rho}$, $m_{\phi}$ and $M_{h}=2M_{N}$, respectively, by $m_{\rho}^{\ast}$, $m_{\phi}^{\ast}$ and $M_{h}^{\ast}=2M_{N}^{\ast}$ in the quark axial current and the radial wave functions of the baryon $B$ are modified by replacing the baryon mass $M_{B}$ by the effective baryon mass in medium $M_{B}^{\ast}$. As for the meson cloud contributions, they change as a consequence of the modification of the coupling $g_{MBB^{\prime}}$ to $g_{MBB^{\prime}}^{\ast}$.
$\rho/\rho_{0}$ $M_{N}^{\ast}$ $m_{\rho}^{\ast}$ $g_{\pi NN}^{\ast}/g_{\pi NN}$ $g_{A}^{N\ast}/g_{A}^{N}$ $G_{A}^{N}(0)$ $n_{P}$
@@
In addition to the bare contributions ($\tilde{G}_{A}^{\rm B}$ replaced by $\tilde{G}_{A}^{{\rm B}*}$), the effect of the nuclear medium acts in two levels: on the meson cloud contribution ($\tilde{G}_{A}^{\rm MC}$ replaced by $\tilde{G}_{A}^{{\rm MC}*}$) and on the wave function normalization factors $Z_{B}$ and $Z_{B^{\prime}}$ (global modification of $\sqrt{Z_{B}Z_{B^{\prime}}}$ to $\sqrt{Z_{B}^{\ast}Z_{B^{\prime}}^{\ast}}$).
We discuss the modifications on the function $\tilde{G}_{A}^{{\rm MC}}$ first. From the discussion from Eq. (29) relative to the meson cloud contribution to a generic $B\to B^{\prime}$ transition, we conclude that $\tilde{G}_{A}^{\rm MC}$ include linear combinations of $F^{\prime}$ and $D^{\prime}$, which are proportional to the product of two coupling constants $g_{\pi BB^{\prime}}^{2}$, since these include diagrams with two pion-baryon couplings (see Fig. 1). The combination of coupling constants can be expressed in terms of $g_{\pi NN}^{2}$, as discussed already. The conclusion is then that $\tilde{G}_{A}^{{\rm MC}}\propto g_{\pi NN}^{2}$ and we can write
@@
In the extension to the nuclear medium we assume, for simplicity, that medium effects associated with the meson cloud contributions are taken into account using the modified coupling constants and normalization factors. Thus, in medium $\tilde{G}_{A,N}^{\rm MC}$ is determined by Eq. (28) where the cutoff $\Lambda$ value is determined by nucleon data in vacuum.
@@
The axial-vector form factors of baryon octet in vacuum were already calculated in Ref. [7]. We consider now the extension to the nuclear medium. An important property observed in the nuclear medium is the suppression of the axial-vector form factors due to the enhancement of the $P$-state component, in general associated with the lower component of the quark Dirac spinors. In the present case the simplest way to include the $P$-state effect is to modify the value of $n_{P}$. In the absence of a solid, established reference, we adjust the value of $n_{P}$ in order to reproduce the suppression estimated by the bag model from Ref. [66]. Since our bare estimate for $G_{A}^{\rm B}(0)$ for the nucleon is similar to the Bag Model $g_{A}^{N}$, we use the in-medium to free space bare axial-vector coupling ratio $g_{A}^{N\ast}/g_{A}^{N}$ to fix the amount of $P$-state admixture (coefficient $n_{P}$). Notice that the determination of $n_{P}$ by the values of $g_{A}^{N\ast}/g_{A}^{N}$ is not an additional constraint of the model. It is a necessary condition associated with the consistency with the QMC model. The QMC model uses the input of the bag model for $g_{A}^{N\ast}$ [66] and this result is used in the calculation of coupling constants $g^{\ast}_{MB_{1}B_{2}}$ [28]. When we include the effect of the meson cloud for the nucleon on the functions $G_{A}^{\ast}$ we modify the values of $G_{A}(0)$ and $G_{A}^{\ast}(0)$, and, consequently, the ratio $G_{A}^{\ast}(0)/G_{A}(0)$ deviates from the $g_{A}^{N\ast}/g_{A}^{N}$.
The reduction of the axial coupling $g_{A}^{N}$ in the nuclear medium is also observed in calculations based on the soliton model [69, 71, 75].
$\rho/\rho_{0}$ $M_{N}^{\ast}$ $M_{\Lambda}^{\ast}$ $M_{\Sigma}^{\ast}$ $M_{\Xi}^{\ast}$ $m_{K}^{\ast}$ $m_{\rho}^{\ast}$ $m_{\phi}^{\ast}$ $g_{\pi NN}^{\ast}/g_{\pi NN}$ $n_{P}$
0 0.9390 1.1160 1.1920 1.3180 0.4937 0.7545 1.0195 1.000 $-0.507$
0.5 0.8313 1.0439 1.1214 1.2822 0.4573 0.7061 1.0191 0.921 $-0.536$
1.0 0.7545 0.9927 1.0704 1.2567 0.4305 0.6537 1.0189 0.899 $-0.560$
Table 6: Model parameters used in the extension to the nuclear medium (baryon octet case). The discussion about the values of $n_{P}$ are in the previous sections. The kaon mass is used in the calculation of the $G_{P}$ pole term in the $|\Delta S|=1$ transitions. For $m_{\pi}^{*}\simeq m_{\pi}$ we use $m_{\pi}=0.138$ GeV [17, 18, 19, 20].
The determination of the value of $n_{P}$ is based on the relation [7]
$\tilde{G}_{A}^{\rm B}(0)=\frac{15-19n_{P}^{2}}{9}g_{A}^{q}(0),$ (37)
where $g_{A}^{q}(0)=1$.
V Numerical results
In this section we present our numerical results for the octet baryon axial form factors $G_{A}$ and $G_{P}$ in vacuum and in medium. In the calculations we use the parameters from Table 6.
As discussed already, the axial transitions between octet baryon members correspond to six $|\Delta I|=1$ transitions (four independent) and six $|\Delta S|=1$ transitions. To simplify the analysis and the discussion of the results, we consider here only typical cases, for the $|\Delta I|=1$ and $|\Delta S|=1$ channels, associated with different octet baryon masses. The comparison of all channels is presented in Appendixes B and C.
We start with the discussion of the results for the vacuum and some remarks about the calibration of our model in vacuum (Sec. V.1). Next we discuss the results in the nuclear medium (Sec. V.2). We display the vacuum and in-medium results side by side in order to better observe the modifications due to the nuclear medium (Figs. 3, 4, 5 and 6).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 3: Axial-vector form factor $G_{A}$ for $|\Delta I|=1$ transitions. Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$). The results for the $\Sigma^{+}\to\Lambda$ and $\Sigma^{0}\to\Sigma^{+}$ transitions are presented in Appendix B, Fig. B1.